In solving -3y+14=98, Paula instructed her CAS to divide both sides by -3.

a. What result did she get?

Answer by edjones(8007)

-3y+14=98
y-(14/3)= -98/3
y=14/3-98/3 add 14/3 to both sides
y= -84/3
y= -28
Check:
-28*-3+14=98
84+14=98 true
Ed